A package contains 10 pens, three of which are defective. If two pens are selected at random, find the probability of getting one defective pen?
Answer by stanbon(75887)

A package contains 10 pens, three of which are defective. If two pens are selected at random, find the probability of getting one defective pen?
------------
P(defective) = 3/10
---
P(not-defective) = 7/10
---
Binomial Problem with n = 2 and p = 0.3
P(x = 1) = 2C1(0.3)(0.7) = 2*0.21 = 0.42
===========================================
Cheers,
Stan H.
==============